Required information Camp Rainbow offers overnight summer camp programs for children ages 10-14 every summer during June andRequired: 1. Perform a least-squares regression analysis on Camp Rainbows data. 2. Using the regression output, create a cosRequired: 1. Perform a least-squares regression analysis on Camp Rainbows data. 2. Using the regression output, create a cosRequired: 1. Perform a least-squares regression analysis on Camp Rainbows data. 2. Using the regression output, create a cos
Answer #1
Week No. of campers(X) Camp Cost(Y)(in $) Square of X X*Y(in $)
1                                         94                     8,630                        8,836                    8,11,220
2                                       140                   11,740                      19,600                  16,43,600
3                                       108                     8,830                      11,664                    9,53,640
4                                       120                     9,400                      14,400                  11,28,000
5                                       118                     9,010                      13,924                  10,63,180
6                                       100                     8,610                      10,000                    8,61,000
7                                         90                     8,680                        8,100                    7,81,200
8                                       124                     9,590                      15,376                  11,89,160
Total                                       894                   74,490                   1,01,900                  84,31,000
Unit Variable Cost (b) = nΣxy − (Σx)(Σy) /nΣx^2 − (Σx)^2
Total Fixed Cost (a) = Σy − bΣx / n
So here n =8 ,Σx =894 Σy =74,490, Σx^2=101,900 ,Σxy =8,431,000
Variable cost per unit(b) =8*8,431,000 - (894*74,490) / 8*101,900 - 894^2
Variable cost per unit(b) =(67,448,000 - 66,594,060) / 815,200 - 799,236
Variable cost per unit(b) = 853,940 / 15,964 =$53.49 per unit
Total Fixed cost =74,490 - ($53.94*894) / 8
Total Fixed cost =(74,490 - 48,222) / 8
Total Fixed cost =$26,268 /8 =$3,283.46
The cost volume formula is:
y =$3,283.46 + $53.49X
Intercept $3,283.46
X Variable 1 $53.49
Total Cost if 90 children attended the session =$3,283.46 + $53.49*90
Total Cost if 90 children attended the session =$3,283.46 + 4,814.10 =$8,097.56